OpenStudy (anonymous):

Suppose g(4) = 4, g'(4) = −2 and f(x) =g(x)/x. Find f'(4).

OpenStudy (mathmate):

From f(x)=g(x)/x, find g'(x). f'(x) =(d/dx)(g(x)*x^-1) =g'(x)/x+g(x)(-x^-2) =g'(x)/x - g(x)/x^2 so f'(4)=g'(4)/4-g(4)/4^2 =-2/4-4/16 =-3/4 Check my arithmetic, my eyes are barely open.